# Alexandra Ciobotaru

### v2

@@ 6 Conclusion and Future Work
PsihoRo is the first open-source dataset for mental health in Romanian. It was created by utilizing a compound form containing both open ended questions for text gathering and the PHQ-9 and GAD-7 consecrated questionnaires to asses the risk of depression and/or anxiety of the respondents. Even if the dataset has only 205 entries, some very interesting insights can be observed.
Regarding the psychologically meaningful categories from LIWC, we observed differences in categories related to body-related words, which may indicate somatic awareness, and tentative words, which indicate uncertainty, that correlated positively with both depression and anxiety scores. A higher preoccupation with mortality was also observed in relation to depression, while anxiety scores were positively correlated with a preoccupation with family. While LIWC analyses showed that lexical indicators can effectively capture underlying psychological patterns related to depression and anxiety, we did not find any association between the linguistic marker ”I” and depression or anxiety in Romanian language, consistent with previous research.
The emotion analysis showed the prevalence of sadness emotions in the population group with risk of depression, as well as fear for the risk of anxiety population group, which is also in concordance with previous research.
The inability to develop a regression model with consistent results based on PsihoRo can be attributed both to the low number of samples and also to the naivety of assuming that depression and anxiety can be reliably discerned from texts, even when supplemented by self-assessment ratings like PHQ-9 and GAD-7. Depression and anxiety require meticulous analysis from multiple perspectives, rather than solely relying on an individual’s textual expression. However, by offering high-quality data with labels for anxiety and depression, PsihoRo seeks to support this intricate analysis.
@@ 7 Ethical Considerations and Limitations
The data collection was approved by The Ethics Committee of the University of Bucharest. All respondents provided their informed consent before answering the survey, which was completely anonymous, no personal data was requested or stored (such as name, surname, email address, IP, etc.).
Adding more data by increasing the number of respondents, by using crowdfunding or similar platforms, could greatly improve PsihoRo and we aim to do this in future versions of this resource.
@@ Acknowledgments
This research was supported by the University of Bucharest and Druid AI.

### v3

@@ 6 Conclusion and Future Work
In this paper, we presented PsihoRo, the first open-source dataset for mental health in Romanian. The dataset was collected from 205 participants who filled in a comprehensive form that included both open-ended questions for text collection and validated questionnaires, specifically the PHQ-9 and GAD-7, to evaluate their risk of depression and anxiety.
We analyzed the collected data using the psychologically meaningful categories from LIWC, emotion analysis and topic modeling. We observed differences in LIWC categories related to body-related words, which may indicate somatic awareness, and tentative words, which indicate uncertainty, that correlated positively with both depression and anxiety scores. A higher preoccupation with mortality was also observed in relation to depression, while anxiety scores were positively correlated with a preoccupation with family. While LIWC analyses showed that lexical indicators can effectively capture underlying psychological patterns related to depression and anxiety, we did not find any association between the linguistic marker “I” and depression or anxiety in the Romanian language, consistent with previous research [36]. The emotion analysis revealed a prevalence of sadness in the population group at risk for depression, as well as fear in the population group at risk for anxiety, which aligns with previous research [25, 3, 38]. Topic modeling showed that although some perceptions and emotions are captured in textual responses, a multi-perspective approach that goes beyond self-expressed text is necessary to fully comprehend these patterns. By offering high-quality data with precise labels for anxiety and depression risk, PsihoRo seeks to supplement this intricate research.
@@ 7 Ethical Considerations and Limitations
The data collection received approval from the Ethics Committee of the University of Bucharest. All respondents provided their informed consent before participating in the survey, which was completely anonymous. No personal information, such as names, surnames, email addresses, or IP addresses, was requested or stored.
We believe that increasing the number of respondents through crowdsourcing or similar platforms could significantly enhance PsihoRo. We aim to implement this in future versions of this resource.
@@ Acknowledgments
This work was partially supported by Druid AI, University of Bucharest and the Ministry of Education and Research, CNCS-UEFISCDI, project SIROLA, number PN-IV-P1- PCE-2023-1701, within PNCDI IV, and by the project "Romanian Hub for Artificial Intelligence - HRIA”, Smart Growth, Digitization, and Financial Instruments Program, 2021-2027, MySMIS no. 334906.